OpenStudy (arivillafana):

Helppppp (^: Consider 2 parallel lines m and n. M<5=95 Find the angles measure below. (I'll draw the angle right now after I post the question...) m<8= m<4= m<3=

OpenStudy (arivillafana):



OpenStudy (xapproachesinfinity):

we have angle 5 =95 by complementary angle angle 8=180-95

OpenStudy (xapproachesinfinity):

then by vertical angles 6=8 next is 1=6 by opposite interior angles (i guess that's what its called)

OpenStudy (xapproachesinfinity):

then by complementary angles again 1+4=180

OpenStudy (arivillafana):

I'm so confused lol

OpenStudy (xapproachesinfinity):

lol i went too fast

OpenStudy (xapproachesinfinity):

the number 1 to 8 angles so don't confusee them with other numbers ok

OpenStudy (xapproachesinfinity):

\[\text{by complementary angles :} ~~~ \angle 8+\angle 5=180 \Longrightarrow \\ \angle 8=180-95=85\]

OpenStudy (xapproachesinfinity):

good so far?

OpenStudy (arivillafana):

ummm sure lol

OpenStudy (xapproachesinfinity):

ok now we go to \[\text{by vertical angles:} ~~~\angle 5\cong \angle 7 \\ \Longrightarrow \angle 7=95\]

OpenStudy (xapproachesinfinity):

got that?

OpenStudy (arivillafana):

I dont mean to sound harsh, but Im in a hurry. I am trying to get my math grade up and was hoping to get the straight-forward answers..

OpenStudy (xapproachesinfinity):

the im sorry can't give you straight forward answer!!

OpenStudy (arivillafana):

ughhh this is so difficult to understand, ill just guess